OpenStudy (anonymous):

Interest Problem: Julian deposits money into two different savings accounts. He deposits $600 more into Account B than Account A. After a certain amount of time, Account A has earned $180 at a rate of 4%, and Account B has earned $375 at a rate of 5%. How much did Julian initially deposit into each account? a.What equation is used to solve this problem? What does each variable represent? t=I/Pr, t-time, I-interest, P-principal, r-rate b.Identify the given information. What do you know about P? 180/(P)0.04=375/(P+600)0.05 Cross Multiply: 375(P)(0.04)=180(P+600)(0.05)

OpenStudy (anonymous):

I already did this part. need help with rest.

OpenStudy (anonymous):

c.Enter the given information into the equation. d.Solve the equation. e.State the solution.

OpenStudy (anonymous):

What does the rest mean?

jimthompson5910 (jim_thompson5910):

Let x = amount deposited in account A and y = amount of time that goes by Account A: I = Prt 180 = x*0.04*y 180/0.04 = xy 4500 = xy y = 4500/x ---------------------------- Account B: I = Prt 375 = (x+600)*0.05*y ... Since he deposits $600 more in this account 375 = (x+600)*0.05*(4500/x) 375x = (x+600)*0.05*4500 375x = 22.5x + 135000 Now solve for x to find the amount deposited in account A (you can then use this to find the amount deposited in account B)

OpenStudy (anonymous):

So is this Part C?

OpenStudy (anonymous):

I have no idea how to solve this!

jimthompson5910 (jim_thompson5910):

375x = 22.5x + 135000 375x - 22.5x = 22.5x + 135000 - 22.5x 352.5x = 135000 352.5x/352.5 = 135000/352.5 x = 382.98 So he deposited about $382.98 in account A and $982.98 in account B (add $600 to the amount in account A to get the amount in account B)

OpenStudy (anonymous):

Ok, thx so much! :) LOL! I don't understand this lesson! :D

jimthompson5910 (jim_thompson5910):

np